OpenStudy (anonymous):

Find all solutions to the equation. cos2x + 2 cos x + 1 = 0

OpenStudy (akashdeepdeb):

When you have trigonometric functions with 'different-angles', try to bring them down to the same angle.\[cos 2x = 2cos^2 x - 1\] So you can write your original equation as: \(2cos^2 x - 1 + 2 cos x + 1 = 0\) Take \(cos x = y\) And then solve the quadratic equation you get. :)

OpenStudy (anonymous):

I don't get it, i thought we're suppose to factor?

OpenStudy (akashdeepdeb):

Yes, you are supposed to factor! After taking cos x = y What do you think the equation becomes? :)

jhonyy9 (jhonyy9):

i think cos2x wann being here cos^2 x or not ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so is it (2cos+1)(cos-1) ? when factored ?

jhonyy9 (jhonyy9):

yes what ?

jhonyy9 (jhonyy9):

what i have supposed ?

OpenStudy (anonymous):

yes it's suppose to be cos^2 x

jhonyy9 (jhonyy9):

cos^2 x +2cos x +1 =0 cos x = t t^2 +2t +1 =0 (t+1)^2 =0 t+1=0 t = -1 cos x = -1 can you continue it ?

OpenStudy (anonymous):

cos x= pi ?

jhonyy9 (jhonyy9):

not cos x = -1 so then x = arccos (-1) do you know how many degree cosinus is equal -1 ?

OpenStudy (anonymous):

180

OpenStudy (anonymous):

\[1+\cos 2x+2\cos x=0\] \[2 \cos ^2 x+2 \cos x=0\] \[2 \cos x \left( \cos x+1 \right)=0\] \[Either ~\cos x=0,~or~\cos x=-1\] solve the two and get the solution.

OpenStudy (anonymous):

\[\cos x=0=\cos \left( 2 n+1 \right)\frac{ \pi }{ 2 },n~ is~ any~ integer.\] \[\cos x=-1=\cos \left( 2 n+1 \right)\pi,where ~n~is~any~integer.\]